# Unified Software Engineering agent as AI Software Engineer

### v1

@@ 4.1. Instantiating AutoCodeRover as USEagent
In this section, we highlight steps to build the unified SE agent USEagent that handles multiple types of SE tasks, from an existing SE agent that only focuses on one task type. We build our USEagent on top of the open-source agent AutoCodeRover (Zhang et al., 2024b; Ruan et al., 2024), since AutoCodeRover operates in different phases which are suitable as starting points for actions. We elaborate the changes to AutoCodeRover and flesh out our proposed solutions to the challenges in Section 4. Fine grained implementation details, prompts, etc. can be found in the supplementary material.
AutoCodeRover is an LLM agent that targets program maintenance tasks. Figure 2 shows the overall workflow of AutoCodeRover. Starting from a natural-language description of a software issue, AutoCodeRover first writes an executable self-contained test that reproduces the issue (Step ①). The stacktrace of the reproducer test, together with the task description, is sent to a context retrieval component to identify relevant code locations (Step ②). The relevant program locations are made available to the Edit Code component for writing a candidate patch (Step ④). With both LLM-generated patch and reproducer test, the Review Patch component then executes the tests on the patched program, decides whether the are correctly written, and iteratively improves them (Step ⑤). Finally, the AutoCodeRover workflow finishes when the reviewer approves a patch or the execution reaches a pre-defined round limit (Step ⑥). AutoCodeRover proved effective for resolving software issues such as bug fixing (Zhang et al., 2024b; Ruan et al., 2024). However, these fixed transitions among components make agents like AutoCodeRover not generalizable to task types that require a different workflow. For example, in a task like test-suite generation, steps such as Reproduction and Review Patch are irrelevant, requiring the design of a new workflow or even a new agent.
@@
Instead of designing a new agent for each task type, our goal is to build one general agent that handles multiple task types. To achieve this, we disassemble the workflow of AutoCodeRover and reassemble it into USEagent shown in Figure 3. We adapt components in AutoCodeRover such as Code Retrieval into actions that can be freely composed by a Meta-Agent for different task types. Given the task description that mentions the task type in natural language, the Meta-Agent uses ReAct-style reasoning to select next actions to execute. Short-term feedbacks such as action execution results are reflected directly to the Meta-Agent, while artifacts that should be preserved longer are stored into a task state representing the consensus memory among actions. We next discuss each component of USEagent in more details.
@@
ExecuteTests Execute a subset of the program test suite List[TestFile] $\rightarrow$ Summary(Results) Read: LocationsTest
@@
EditCode Make modifications across the codebase, including test-code. $\Delta$Behavior $\times$ Location $\rightarrow$ Diff Read: LocationsCode
@@
ReviewPatch Review and iteratively improve a generated patch and generated test. TestReprod.$\times$×DiffCode $\rightarrow$ Bool Read: DiffCode
@@
Given the tasks of program repair, regression testing, code generation, and test generation in USEbench, we identify a priming set of actions for USEagent. The set of actions are shown in Table 2. Each action exposes its interface consisting of the description, input and output to the Meta-Agent, but does not expose execution details within the action. In addition to the input/output interaction with Meta-Agent, each action can also read/write to the task state.
Many of the actions in Table 2 come from existing agents like AutoCodeRover (Figure 2), and incorporating them into USEagent only involves specifying their input/output interface. On top of the existing agents, we designed a few new actions that are essential for the task types in USEbench. These new actions include TestRetrieval and ExecuteTests. The TestRetrieval action explores the codebase and retrieves testcases relevant to the current task. It works similarly compared to the existing CodeRetrieval, which employs a set of search tools such as search_func to search for relevant code units in the program. The ExecuteTests action executes parts of the project test suite, which is useful for validating code/test changes by execution. Current agents focusing on SWE-bench usually assume the commands to run existing unit tests are given as an input to the agent (Ruan et al., 2024; Xia et al., 2024); however, this assumption requires additional manual configuration when setting up the agents on new projects. Our ExecuteTests action does not assume the test commands to be given, and instead it queries the project documentation files to retrieve the project-specific commands through RAG. In addition, we added a special Finish action that signals the termination of the agent execution. The Finish action is invoked with an argument specifying the final result for the current task. Unlike other actions which encapsulates concrete workflows, the Finish action only terminates the agent and outputs its argument as the final result from the agent.
@@
Our next contribution towards a USEagent is the design of a task state. The task state represents the consensus memory among the actions - each action can write new artifacts generated by them into the task state so that the artifacts can be visible to other actions. Our task state captures essential artifacts involved in a software debugging process, including relevant locations, test execution information, and prior attempts of code modifications. Specifically, the task state $S$ is defined as: $S=(L_{c},L_{t},R_{exec},DS)$. Here, $L_{c}$ represents the relevant code locations (e.g., program methods and classes); $L_{t}$ represents the test locations (e.g., unit test methods). The locations are identified by the retrieval actions, and are utilized by downstream code editing actions. $R_{exec}$ contains the execution results of different tests against different candidate patches. These execution information provides guidance to other actions in improving partial solutions and selecting the final solution. The last part of the task state is a diff store ($DS$), which records all the generated diff contents from the EditCode, Reproduction and ReviewPatch actions. These diff contents include modifications to both the program code and the test code. The diff store is necessary to allow for versatile invocations of the ExecuteTests and EditCode actions. By selecting different subsets of diffs in $DS$ and passing them to ExecuteTests, tests can be executed on different versions of the project after applying the selected diffs. Similarly, EditCode can first apply an existing patch from $DS$, and subsequently introduce additional modifications, thereby enhancing partial patches.
@@ 4.2. Configuring OpenHands
OpenHands CodeActAgent (Wang et al., 2024b) is a general-purpose agent designed to solve general tasks [^1] Note: OpenHands is also developing CodeActSWEAgent, specialised on SWE-Bench. We purposefully chose the non-specialised implementation for USEbench. The entries on the SWE-Bench leaderboards are to our knowledge results from CodeActSWEAgent.. It utilizes a structured controller-agent-runtime architecture, where the AgentController serves as a supervisor that enforces operational constraints (conversation iterations, budget) and manages the agent’s lifecycle (start/stop/pause), while the CodeActAgent itself makes decisions, interpreting LLM responses and converting them into actions to interact with the runtime, which are isolated sandbox environment. Since CodeActAgent does not rely on a predefined task-specific workflow, it can be applied to a wide range of software engineering tasks, including those in USEbench, without requiring modifications to its architecture.
@@
The execution of CodeActAgent terminates when the agent issues an AgentFinishAction, reaches resource limits (e.g., maximum iterations), or encounters an error. In practice, we design different prompt templates for each task type as input to OpenHands CodeActAgent, incorporating both the task type and task description. Compared to USEagent, the OpenHands CodeActAgent employs a similar reasoning framework, but the actions operate on a lower-level such as execution of a single bash command. It follows a single ReAct-loop centered around a more open set of available actions. This free-wheeling approach is an alternative to the more structured and pronounced options in Section 4.1.
@@ 5. Research Questions & Experiment Setup
RQ1: Efficacy of Agentic Systems How well do state of the art agentic systems work when facing varying software engineering tasks from USEbench?
We answer RQ1 by adapting AutoCodeRover and OpenHands before applying them to all data points outlined in Section 2 and report a PASS@1. In addition to the efficacy, we also report an investigation of the major obstacles for the individual benchmarks. To gather information on retrys and estimate the effect of randomness, we perform a PASS@5 on a significant subset of the datapoints.
A known issue and motivation for our next research question is overfitting in plausible patches. Such patches pass the evaluation-criteria by bypassing the harness with tricks rather than benign functionality. Sometimes, actual solutions are checks for edge-cases - these can only be distinguished from overfitting by consulting the ground truth (i.e., the gold patch). We aim to identify (and quantify) false-positives and other cases of overfitting solutions.
RQ2: Analysis of Plausible Patches What is the rate of false positives among the resolved patches?
We sample a significant subset222at 95% confidence with 10% error margin of solved datapoints and manually examine whether the plausible solution is overfitting. During this manual investigation, we also look for memorization in plausible patches, i.e. patches that perfectly mirror the ground truth developer patch or contain suspicious content.
@@
RQ3: Effectiveness of self-configuration Can USEagent self-configure to different task types?
We present the sequence of actions of the USEagent and accumulate their patterns per task. We showcase examples on how these top-level patterns translate into commands, queries and edits.
@@
RQ4: Open challenges for agentic systems What are the challenges faced by the current unified agent in tasks in USEagent? Can we identify actions to remediate them?
@@
We evaluate agentic systems such as USEagent and OpenHands CodeActAgent on the full USEbench which consists of 1271 datapoints. In our experiments, all agentic systems use Anthropic Claude 3.5 Sonnet v2 (claude-3-5-sonnet-20241022) as the backend LLM. For USEagent, we set the maximum rounds of action invocation from the Meta-Agent to 20, and force the agent to end execution and select a candidate solution. We set the model temperature to zero in USEagent. To better understand the effect of randomness, we randomly sampled a statistically significant subset (295 instances) to report PASS@5 results from both agentic systems.
@@ 6.1. RQ1: Efficacy of Agentic Systems
Dataset Tasks PASS@1 PASS@5 PASS@1 PASS@5
SWE-Ver 500 228 (45.6%) 66.7% 192 (38.4%) 52.1%
SWT 298 120 (40.3%) 53.6% 85 (28.4%) 50.7%
REPOCOD 200 12 (6.0%) 15.2% 11 (5.5%) 6.5%
REPOTEST 173 55 (31.8%) 42.5% 45 (26.0%) 42.5%
SWETRY 100 8 (8.0%) 30.4% 7 (7.0%) 8.7%
Total 1271 423 (33.3%) 49.5% 340 (26.8%) 44.1%
@@
[…]
@@
While being effective on issue resolution tasks, USEagent can handle other tasks which specialized agents could not be applied to. For issue reproduction tasks in SWT, USEagent resolved 40.3% of the tasks. Here, “resolved” means USEagent generated test cases that cover all changed lines in the developer-written patch for the issue, without seeing the patch. Moreover, for a significant number of unresolved tasks, USEagent generated test cases that achieve a coverage above 90%. In these cases, the generated tests do not fully satisfy the requirement of the benchmark, but only miss coverage on individual statements or branches. On test generation tasks in REPOTEST, USEagent achieves efficacy of 31.8%, which is slightly lower than similar tasks in SWT. While the requirement in SWT is to cover a developer-written patch, the requirement in REPOTEST is to generate tests that cover an entire method, which can be more challenging. There are generated tests that cover most scenarios in a method, but do not cover all the lines. For example, one pattern in the unresolved tasks is that the agent provides only one positive test-input that executes the “normal” paths of a method, followed by tests that raise exceptions. In contrast, the ground-truth tests written by developers not only focus on exception paths, but extensively test the “normal” inputs. We argue for certain advantages in agent-generated tests over the developer-written tests; the agent often generates shorter tests where each test has individual assertions, while the developer-written tests tend to consist of long tests that combine multiple inputs and assertions. In a real test-suite, short individual tests can provide more informative error messages when part of the test-suite fail (Spadini et al., 2018; Van Deursen et al., 2001).
@@
Tasks in REPOCOD proved to be the most challenging with only 6% resolution rate. REPOCOD tasks are generally challenging because the requirement is to generate a complete method implementation that can pass a large number of hidden test cases. Overall, for many unresolved instances, the generated method passes a high number of tests, yet does not account for a few edge-cases. The majority of resolved instances stem from Sympy, a library for symbolic mathematics. We expect a relation between the relative success in Sympy and the fact that mathematical reasoning is becoming a standard element of training data in foundational models.
@@
On the compound benchmark SWETRY, USEagent has a resolve rate of 8%, which shows some promise to apply USEagent as a follow up to partial fixes. We need to stress that the datapoints in SWETRY are sourced from previous failed attempts of agentic systems, which means they are from the more challenging instances in SWE-verified. We observe two main failure modes on the SWETRY tasks. Firstly, since the partial patch is given to the agent as part of the task description, the agent often attempts to derive a solution on top of this partial patches by making modifications to it. However, the partial patch may be misplaced. In these cases, the agent still chose to iterate over this patch, and failed to discard the partial attempt and approach the task in a different way. Secondly, the Meta-Agent sometimes ends the agent execution prematurely. For example, when the ExecuteTests action still shows some relevant test errors, the Meta-Agent wrongly disregards errors as irrelevant to the task, ending the workflow. This can be addressed by configuring the prompt, at the cost of longer execution time and potential overfitting in other benchmarks.
@@
OpenHands CodeActAgent achieves a PASS@1 of 38.4% on SWE-verified, which is lower than the reported result from the SWE-Bench leaderboard (SWE, [n. d.]). We attribute this difference to the choice of agent - the reported result on the leaderboard is generated using a specialized CodeActSWEAgent, while we used the more general CodeActAgent in our evaluation.
@@
Moving to a PASS@5 increases the efficacy of USEagent to 49.5% (+48.6%) and OpenHands to 44.1% (+64.6%). This shows a similar increase across the two systems when employing global retries. Both systems largely benefit when facing edge-cases, e.g. to cover the last statement in testing or finding a correct reproducer.
Summary: RQ1 - Efficacy of Agentic Systems USEagent solves the tasks from USEbench at a 33.3% rate, compared to 26.8% of OpenHands. There is a consistent improvement over all benchmarks. USEagent’s efficacy of 45.6% for SWE-verified is close to the performance of fine-tuned systems.

### v2

@@ 4.1. Instantiating AutoCodeRover as USEagent
In this section, we highlight steps to build the unified SE agent USEagent that handles multiple types of SE tasks. We design our USEagent drawing inspiration from the open-source agent AutoCodeRover (zhang2024autocoderover; ruan2024specrover), since AutoCodeRover operates in different phases which are suitable as starting points for actions. We elaborate the design of USEagent and flesh out our proposed solutions to the challenges in Section 4.
AutoCodeRover is an LLM agent that targets program maintenance tasks. Figure 2 shows the overall workflow of AutoCodeRover. Starting from a natural-language description of a software issue, AutoCodeRover first writes an executable self-contained test that reproduces the issue (Step ①). The stacktrace of the reproducer test, together with the task description, is sent to a context retrieval component to identify relevant code locations (Step ②). The relevant program locations are made available to the Edit Code component for writing a candidate patch (Step ④). With both LLM-generated patch and reproducer test, the Review Patch component then executes the tests on the patched program, decides whether the are correctly written, and iteratively improves them (Step ⑤). Finally, the AutoCodeRover workflow finishes when the reviewer approves a patch or the execution reaches a pre-defined round limit (Step ⑥). AutoCodeRover proved effective for resolving software issues such as bug fixing (zhang2024autocoderover; ruan2024specrover). However, these fixed transitions among components make agents like AutoCodeRover not generalizable to task types that require a different workflow. For example, in a task like test-suite generation, steps such as Reproduction and Review Patch are irrelevant, requiring the design of a new workflow or even a new agent.
@@
Instead of designing a new agent for each task type, our goal is to build one general agent that handles multiple task types. To achieve this, we disassemble the workflow of AutoCodeRover and reassemble it into USEagent shown in Figure 3. We adapt components in AutoCodeRover such as Code Retrieval into actions that can be freely composed by a Meta-Agent for different task types. Given the task description that mentions the task type in natural language, the Meta-Agent uses ReAct-style reasoning to select next actions to execute. Short-term feedback such as action execution results are reflected directly to the Meta-Agent, while artifacts that should be preserved longer are stored into a task state representing the consensus memory among actions. We next discuss each component of USEagent in more details.
@@
ExecuteTests Execute a subset of the program test suite. List[TestFile] $\rightarrow$ Summary(Results) Read: LocationsTest
@@
[…]
@@
[…]
@@
Given the tasks of program repair, regression testing, code generation, and test generation in USEbench, we identify a priming set of actions for USEagent. The set of actions are shown in Table 2. Each action provides the Meta-Agent with an interface that specifies its description, input, and output, but abstracts away the underlying execution details. In addition to interacting with the Meta-Agent through input and output, each action can also read from and write to the task state.
Many of the actions in Table 2 are conceptually inspired by those used in existing agents such as AutoCodeRover (Figure 2). On top of the existing agents, we designed a few new actions that are essential for the task types in USEbench. These new actions include TestRetrieval and ExecuteTests. The TestRetrieval action explores the codebase and retrieves testcases relevant to the current task. It works similarly compared to the existing CodeRetrieval, which employs a set of search tools such as search_func to search for relevant code units in the program. The ExecuteTests action executes parts of the project test suite, which is useful for validating code/test changes by execution. Current agents focusing on SWE-bench usually assume the commands to run existing unit tests are given as an input to the agent (ruan2024specrover; xia2024agentless); however, this assumption requires additional manual configuration when setting up the agents on new projects. Our ExecuteTests action does not assume the test commands to be given, and instead it queries the project documentation files to retrieve the project-specific commands through RAG. In addition, we added a special Finish action that signals the termination of the agent execution. The Finish action is invoked with an argument specifying the final result for the current task. Unlike other actions which encapsulates concrete workflows, the Finish action only terminates the agent and outputs its argument as the final result from the agent.
@@
Our next contribution towards a USEagent is the design of a task state. The task state represents the consensus memory among the actions - each action can write new artifacts generated by them into the task state so that the artifacts can be visible to other actions. Our task state captures essential artifacts involved in a software debugging process, including relevant locations, test execution information, and prior attempts of code modifications. Specifically, the task state $S$ is defined as: $S=(L_{c},L_{t},R_{exec},DS)$. Here, $L_{c}$ represents the relevant code locations (e.g., program methods and classes); $L_{t}$ represents the test locations (e.g., unit test methods). The locations are identified by the retrieval actions, and are utilized by downstream code editing actions. $R_{exec}$ contains the execution results of different tests against different candidate patches. These execution information provides guidance to other actions in improving partial solutions and selecting the final solution. The last part of the task state is a diff store ($DS$), which records all the generated diff contents from the EditCode, Reproduction and ReviewPatch actions. These diff contents include modifications to both the program code and the test code. The diff store is necessary to allow for versatile invocations of the ExecuteTests and EditCode actions. By selecting different subsets of diffs in $DS$ and passing them to ExecuteTests, tests can be executed on different versions of the project after applying the selected diffs. Similarly, EditCode can first apply an existing patch from $DS$, and subsequently introduce additional modifications, thereby enhancing partial patches. Concretely, when an action is implemented, the implementation defines which part of the action output will be stored into the task state.
@@ 4.2. Configuring OpenHands
OpenHands CodeActAgent (wang2024openhands) is a general-purpose agent designed to solve general tasks [^1] Note: OpenHands is also developing CodeActSWEAgent, specialised on SWE-Bench. We purposefully chose the non-specialised implementation for USEbench. The entries on the SWE-Bench leaderboards are to our knowledge results from CodeActSWEAgent.. It utilizes a structured controller-agent-runtime architecture, where the AgentController acts as a supervisor, enforcing operational constraints (such as conversation iterations and budget) and managing the agent’s lifecycle (start, stop, pause). The CodeActAgent makes decisions, interprets LLM responses, and converts them into actions to interact with the runtime, which is an isolated sandbox environment. Since CodeActAgent does not rely on a predefined task-specific workflow, it can be applied to a wide range of software engineering tasks, including those in USEbench, without requiring modifications to its architecture.
@@
The execution of CodeActAgent terminates when the agent issues an AgentFinishAction, reaches resource limits (e.g., maximum iterations), or encounters an error. In practice, we design different prompt templates for each task type as input to OpenHands CodeActAgent, incorporating both the task type and task description. Compared to USEagent, the OpenHands CodeActAgent employs a similar reasoning framework, but the actions operate on a lower-level such as execution of a single bash command. It follows a single ReAct-loop centered around a more open set of available actions. This free-wheeling approach is an alternative to the more structured and pronounced options in Section 4.1. These two designs demonstrate a trade-off: allowing arbitrary bash commands makes the agent more flexible, while using structured actions improves interpretability and control over the agent.
@@ 5. Research Questions & Experiment Setup
RQ1: Efficacy of Agentic Systems.  How well do state-of-the-art agentic systems perform across the diverse software engineering tasks in USEbench?
We address RQ1 by applying USEagent and OpenHands to all data points outlined in Section 2 and report a PASS@1. In addition to the efficacy, we also report an investigation of the major obstacles for the individual benchmarks. To understand the effect of randomness, we report PASS@5 on a significant subset of the data points. We also examine the feasibility of using open-source models as the backend of USEagent, by evaluating USEagent with DeepSeek-V3.
A known issue and motivation for our next research question is overfitting in plausible patches. Such patches pass the evaluation-criteria by bypassing the harness with tricks rather than benign functionality. Sometimes, actual solutions are checks for edge-cases - these can only be distinguished from overfitting by consulting the ground truth (i.e., the gold patch). We aim to identify and quantify overfitting solutions.
RQ2: Analysis of Plausible Patches.  What is the rate of overfitting among the resolved patches?
We sample a significant subset222at 95% confidence with 10% error margin. of solved data points and manually examine whether the plausible solution is overfitting. During this manual investigation, we also look for memorization in plausible patches, i.e. patches that perfectly mirror the ground truth developer patch or contain suspicious content.
@@
RQ3: Effectiveness of Self-Configuration and Actions.  Can USEagent self-configure based on the task type while maintaining reasonable efficacy? How does adding new actions affect efficacy?
We present the sequence of actions of the USEagent and accumulate their patterns per task. We showcase examples on how these top-level patterns translate into commands, queries and edits. Furthermore, to quantitatively understand the effect of using dynamic decision making, we perform an ablation study by using a static action sequence for each task type. We also perform an ablation study by removing one of the newly added actions ExecuteTests.
@@
RQ4: Open Challenges for Agentic Systems.  What are the challenges faced by the current unified agent in tasks in USEagent? Can we identify actions to remediate them?
@@
We evaluate agentic systems such as USEagent and OpenHands CodeActAgent on the full USEbench which consists of 1271 datapoints. In most of our experiments, agentic systems use Anthropic Claude 3.5 Sonnet v2 (claude-3-5-sonnet-20241022) as the backend LLM. Additional experiments with DeepSeek-V3 invokes the DeepSeek-V3 model through the Openrouter API  [^3] https://openrouter.ai/deepseek/deepseek-chat-v3-0324. For USEagent, we set the maximum rounds of action invocation from the Meta-Agent to 20, and force the agent to end execution and select a candidate solution. We set the model temperature to zero in USEagent. To better understand the effect of randomness, we randomly sampled a statistically significant subset (295 instances) to report PASS@5 results from both agentic systems.
@@ 6.1. RQ1: Efficacy of Agentic Systems
Dataset Tasks [Uncaptioned image]PASS@1 [Uncaptioned image] [^4] Derived from the statistically significant subset. PASS@5 [Uncaptioned image] [^4] Derived from the statistically significant subset. PASS@1 [Uncaptioned image]PASS@1 [Uncaptioned image] [^4] Derived from the statistically significant subset. PASS@5 [Uncaptioned image] [^4] Derived from the statistically significant subset. PASS@1
SWE-Ver 500 228 (45.6%) 66.7% 42.7% 192 (38.4%) 52.1% 35.9%
SWT 298 120 (40.3%) 53.6% 34.8% 85 (28.4%) 50.7% 29.0%
REPOCOD 200 12 (6.0%) 15.2% 6.5% 11 (5.5%) 6.5% 0%
REPOTEST 173 55 (31.8%) 42.5% 17.5% 45 (26.0%) 42.5% 10.0%
SWETRY 100 8 (8.0%) 30.4% 8.7% 7 (7.0%) 8.7% 4.3%
Total 1271 423 (33.3%) 49.5% 29.1% 340 (26.8%) 44.1% 22.6%
@@
In general, USEagent demonstrates its applicability across different types of tasks in USEbench. USEagent achieves a PASS@1 of 45.6% on SWE-Verified. In comparison, AutoCodeRover, a more specialized agent on issue resolution, achieved efficacy of 46.2% on the same SWE-Verified benchmark when using the same Claude 3.5 Sonnet v2 backend LLM (SWE-bench-leaderboard). USEagent relaxed the fixed workflow in AutoCodeRover and is a more general agent that applies to more task types. Although being more general, USEagent demonstrates similar efficacy on the specialized issue resolution task compared to AutoCodeRover. This comparable performance demonstrates that, with the design of USEagent, the increased autonomy does not result in noticeable reduction in efficacy.
@@
While being effective on issue resolution tasks, USEagent can handle other tasks which specialized agents could not be applied to. For issue reproduction tasks in SWT, USEagent resolved 40.3% of the tasks. Here, “resolved” means USEagent generated test cases that cover all changed lines in the developer-written patch for the issue, without seeing the patch. Moreover, for a significant number of unresolved tasks, USEagent generated test cases that achieve a coverage above 90%. In these cases, the generated tests do not fully satisfy the requirement of the benchmark, but only miss coverage on individual statements or branches. On test generation tasks in REPOTEST, USEagent achieves efficacy of 31.8%, which is slightly lower than similar tasks in SWT. While the requirement in SWT is to cover a developer-written patch, the requirement in REPOTEST is to generate tests that cover an entire method, which can be more challenging. Sometimes the agent generates tests that cover most scenarios in a method, but do not cover all the lines. On the other hand, we observe certain advantages of agent-generated tests compared to developer-written tests. The agent often generates shorter tests where each test has individual assertions, while the developer-written tests tend to consist of long tests that combine multiple inputs and assertions. In a real test-suite, short individual tests can provide more informative error messages when part of the test-suite fail (spadini2018relation; van2001refactoring).
@@
Tasks in REPOCOD proved to be the most challenging with only 6% resolution rate. REPOCOD tasks are generally challenging because the requirement is to generate a complete method implementation that can pass a large number of hidden test cases. Overall, for many unresolved instances, the generated method passes a high number of tests, yet does not account for a few edge-cases. The majority of resolved instances stem from Sympy, a library for symbolic mathematics. We expect a relation between the relative success in Sympy and the fact that mathematical reasoning is becoming a common aspect of training in foundational models.
@@
On the compound benchmark SWETRY, USEagent has a resolve rate of 8%, which shows some promise to apply USEagent as a follow up to partial fixes. We need to stress that the data points in SWETRY are sourced from previous failed attempts of agentic systems, which means they are from the more challenging instances in SWE-verified. We observe two main failure modes on the SWETRY tasks. Firstly, since the partial patch is given to the agent as part of the task description, the agent often attempts to derive a solution on top of this partial patches by making modifications to it. However, the partial patch may be misplaced. In these cases, the agent still chose to iterate over this patch, and failed to discard the partial attempt and approach the task in a different way. Secondly, the Meta-Agent sometimes ends the agent execution prematurely. For example, when the ExecuteTests action still shows some relevant test errors, the Meta-Agent wrongly disregards errors as irrelevant to the task, ending the workflow. This can be addressed by configuring the prompt, at the cost of longer execution time.
@@
OpenHands CodeActAgent achieves a PASS@1 of 38.4% on SWE-verified, which is lower than the reported result from the SWE-Bench leaderboard (SWE-bench-leaderboard). We attribute this difference to the choice of agent - the reported result on the leaderboard is generated using a specialized CodeActSWEAgent, while we used the more general CodeActAgent in our evaluation.
@@
Open-Source Model
We also report the efficacy of the agents with the open-source model DeepSeek-V3 in Table 3. With DeepSeek-V3, both USEagent and OpenHands have a slight performance drop, which is expected. There is a larger performance drop on REPOTEST for both agents. REPOTEST requires generating a large amount of test code to achieve full test coverage – DeepSeek-V3 sometimes makes minor mistakes in the generated code, or generates tests that only partially cover the target function.
@@
Moving to a PASS@5 increases the efficacy of USEagent to 49.5% (+48.6%) and OpenHands to 44.1% (+64.6%). This shows a similar increase across the two systems when employing retries.
Summary: RQ1 - Efficacy of Agentic Systems.USEagent solves the tasks from USEbench at a 33.3% rate, compared to 26.8% of OpenHands. There is a consistent improvement over all benchmarks. USEagent’s efficacy of 45.6% for SWE-verified is close to the performance of fine-tuned systems.